# Paired Parton Trial States for the Superfluid-Fractional Chern Insulator Transition

### v1

@@ III Comparison to DMRG on the hexagonal model
IV Paired state details
IV.1 Bogoliubov transformation and the mean-field wavefunction form
@@
where $\Psi_{\mathbf{k}}=(f_{1,\mathbf{k}},\ldots,f_{D,\mathbf{k}},f^{\dagger}_{1,\mathbf{-k}},\ldots,f^{\dagger}_{D,-\mathbf{k}})$ and $\Delta(-\mathbf{k})=-\Delta^{T}(\mathbf{k})$. Both $T({\bf k})$ and $\Delta({\bf k})$ are $D\times D$ matrices in this sublattice space. So the matrix in Eq.S1 is $2D\times 2D$ so we get $2D$ BdG bands. From the BdG form, we expect the MF GS $\left.\middle|\Omega\right\rangle$ to fill $D$ of these bands. What follows is a generalization of Read and Green’s logic [26] to multiple lattice sites. Following Read and Green, the $D$ empty bands correspond to operators
@@
for $l=1...D$. The Latin indices represent bands and the Greek indices represent (parton) lattice sites in the model. The ground state is then determined by the constraint that $\alpha_{l\mathbf{k}}\left.\middle|\Omega\right\rangle=0$ for all $\mathbf{k}$ and $l$. It may be written as (prime on the product indicates to only multiply over one of each $\mathbf{k},-\mathbf{k}$ pair)
$\left.\middle|\Omega\right\rangle$ $=\text{const}\prod_{\mathbf{k}}^{\prime}\exp\left[(g_{\mathbf{k}})_{\sigma\rho}f_{\sigma,\mathbf{k}}^{\dagger}f^{\dagger}_{\rho,\mathbf{{-k}}}\right]\left.\middle|\emptyset\right\rangle$ (S3)
@@
$\begin{split}f_{\beta,\mathbf{k}}\left.\middle|\Omega\right\rangle&=\text{const}f_{\beta,\mathbf{k}}\prod_{\mathbf{q}}^{\prime}\exp\left[(g_{\mathbf{q}})_{\sigma\rho}f_{\sigma,\mathbf{q}}^{\dagger}f^{\dagger}_{\rho,\mathbf{{-q}}}\right]\left.\middle|\emptyset\right\rangle=\\ &=\text{const}\left\{\prod_{\mathbf{q}\neq\pm\mathbf{k}}^{\prime}\exp\left[(g_{\mathbf{q}})_{\sigma\rho}f_{\sigma,\mathbf{q}}^{\dagger}f^{\dagger}_{\rho,\mathbf{{-q}}}\right]\right\}f_{\beta,\mathbf{k}}\exp\left[(g_{\mathbf{k}})_{\sigma\rho}f_{\sigma,\mathbf{k}}^{\dagger}f^{\dagger}_{\rho,\mathbf{{-k}}}\right]\left.\middle|\emptyset\right\rangle\\ \end{split}$ (S5)
Then, writing $G=(g_{\mathbf{k}})_{\sigma\rho}f_{\sigma,\mathbf{k}}^{\dagger}f^{\dagger}_{\rho,\mathbf{{-k}}}$, we can see that $[f_{\beta,\mathbf{k}},G]=(g_{\mathbf{k}})_{\beta\rho}f^{\dagger}_{\rho,\mathbf{-k}}$ from which it follows (with $e^{-A}Be^{A}=B+[B,A]+\frac{1}{2!}[[B,A],A]+\frac{1}{3!}[[[B,A],A],A]+\ldots$) that $f_{\beta,\mathbf{k}}e^{G}=e^{G}\left(f_{\beta,\mathbf{k}}+(g_{\mathbf{k}})_{\beta\rho}f^{\dagger}_{\rho,\mathbf{-k}}\right)$.Using that in Eq. S5 along with $f_{\beta,\mathbf{k}}\left.\middle|\emptyset\right\rangle=0$ and $e^{G}f^{\dagger}_{\rho,-\mathbf{k}}=f^{\dagger}_{\rho,-\mathbf{k}}e^{G}$, we get that
@@
IV.2 Recovering mean-field expectations
@@
This is equivalent to the more well-known condition requiring that the Bogoliubovs must obey both $\{\alpha_{l\mathbf{k}},\alpha^{\dagger}_{m\mathbf{q}}\}=\delta_{lm}\delta_{\mathbf{kq}}$ and $\{\alpha_{l\mathbf{k}},\alpha_{m\mathbf{q}}\}=0$. In what follows, we interpret the $2D\times D$ matrix $\begin{pmatrix}U_{\mathbf{k}}^{\dagger}\\ -V_{\mathbf{k}}^{\dagger}\end{pmatrix}$ and similar constructions as a set of $D$ vectors with $2D$ components each. Then, the condition that $U_{\mathbf{k}}g_{\mathbf{k}}=V_{\mathbf{k}}$, is really just the requirement that $\begin{pmatrix}U_{\mathbf{k}}^{\dagger}\\ -V_{\mathbf{k}}^{\dagger}\end{pmatrix}$ is orthogonal to $\begin{pmatrix}g_{\mathbf{k}}\\ \mathbb{1}\end{pmatrix}$, ie that each of the $D$ vectors in one is orthogonal to each of the $D$ in the other. Because $\begin{pmatrix}g_{\mathbf{k}}\\ \mathbb{1}\end{pmatrix}$ is of rank $D$ (guaranteed by the $D\times D$ identity), the dimension of the space orthogonal to it has dimension $2D-D=D$. In fact, it may be checked that a set of $D$ linearly independent vectors orthogonal to it are just the columns of $\begin{pmatrix}\mathbb{1}\\ -g_{\mathbf{k}}^{\dagger}\end{pmatrix}$. So all the vectors of $\begin{pmatrix}U_{\mathbf{k}}^{\dagger}\\ -V_{\mathbf{k}}^{\dagger}\end{pmatrix}$ are within the span of the vectors in $\begin{pmatrix}\mathbb{1}\\ -g_{\mathbf{k}}^{\dagger}\end{pmatrix}$. Furthermore, unitarity of Eq. S8 necessitates among other things $U_{\mathbf{k}}^{\dagger}U_{\mathbf{k}}+V_{\mathbf{k}}^{\dagger}V_{\mathbf{k}}=\mathbb{1}$, in other words the vectors in $\begin{pmatrix}U_{\mathbf{k}}^{\dagger}\\ -V_{\mathbf{k}}^{\dagger}\end{pmatrix}$ are orthonormal. This all implies that we may take
@@
So given a $g(\mathbf{r})$, we have an algorithm which both figures out the appropriate normalization and then computes the BdG Chern number, pairing strength and can also in principle compute any other correlation function at the mean-field level, as Wick’s theorem holds there. Still, we might want to further address the question of why one singles out the particular mean field state where $N(\theta^{*})=NK$. After all, the other $\theta$ which represent different parton mean-field states produce the same projected bosonic wavefunction. The main reason is that the mean-field parton wavefunction lies in a larger Hilbert space, much of which is unphysical — demanding that $N(\theta^{*})=NK$ guarantees that as large a component of that wavefunction is actually in the physical subspace as possible, meaning that mean-field values evaluated at $\theta^{*}$ should be more meaningful for the properties of the physical state — we again note that one has to be careful when interpreting mean-field values physically.
IV.3 Paired Chern numbers and transitions
In the unpaired, $\Delta=0$ and uncorrelated limit, the BdG Chern number is just $\mathcal{C}^{\text{Bdg}}=2(\mathcal{C}^{\text{P}}_{1}+\mathcal{C}^{\text{P}}_{2})$. To see why the factor of $2$, we note that the BdG Chern number counts the number of Majorana edge modes, while the usual Chern numbers count the Dirac edge modes. So we expect $\mathcal{C}^{\text{Bdg}}_{\text{SF}}=0$ and $\mathcal{C}^{\text{Bdg}}_{\text{FCI}}=4$ in the unpaired limit. Assuming a small pairing $\Delta$ does not close any gaps, this should also hold at mean-field for the paired states. We still have the projective translation symmetry for partons $T_{1}^{\text{P}}T_{2}^{\text{P}}=-T_{2}^{\text{P}}T_{1}^{\text{P}}$ which means a doubling in the spectrum, implying that the Chern number jumps by two units, or really double the jump that would happen otherwise. Furthermore, the BdG nature of the problem dictates that a band gap closure at a point $\mathbf{k}$ necessitates another closure at $-\mathbf{k}$. Following logic similar to [13], we may then argue that a generic band-gap closure occurs at four points in the BZ, $\mathbf{k},\mathbf{k}+(0,\pi),\mathbf{-k}$ and $\mathbf{-k}+(0,\pi)$. This leads to the conclusion that the jump from $\mathcal{C}^{\text{Bdg}}=4$ to $\mathcal{C}^{\text{Bdg}}=0$ is natural and that the SF-FCI transition may occur without fine-tuning.
@@ IV.3 Paired Chern numbers and transitions
One way this $\mathcal{C}^{\text{BdG}}=2$ phase could be realized is with parton $f_{2}$ in a $\mathcal{C}=1$ (non-BdG) band, while $f_{1}$ transitions from a $\mathcal{C}^{\text{P}}=1$ (non-BdG) band into a $\mathcal{C}=0$ (BdG) band via a double-Majorana gap closure as discussed above. Following Barkeshli and McGreevy [11], we characterize this phase by its electromagnetic response. The parton $f_{1}$ feels a gauge field $\frac{1}{2}A_{\mu}+a_{\mu}$ while $f_{2}$ feels $\frac{1}{2}A_{\mu}-a_{\mu}$ (here $A_{\mu}$ is external electromag field and $a_{\mu}$ is the emergent $U(1)$ “gluing” gauge field from the parton construction). Since $f_{1}$ is in a paired state, the gauge field it feels, $b_{\mu}=\frac{1}{2}A_{\mu}+a_{\mu}$ gets Higgsed, ie we get a term $cb_{\mu}b^{\mu}\subset\mathcal{L}$. Integrating out the other parton, $f_{2}$ gives the simple IQHE response $\frac{1}{4\pi}\epsilon^{\mu\nu\lambda}(\frac{1}{2}A-a)_{\mu}\partial_{\nu}(\frac{1}{2}A-a)_{\lambda}\subset\mathcal{L}$. Now when integrating out $a_{\mu}$, the most relevant at large distance is the Higgs term which simply sets $b_{\mu}\rightarrow 0$ meaning that we get $a_{\mu}\rightarrow-\frac{1}{2}A_{\mu}$ and so the final electromagnetic response is simply $\mathcal{L}=\frac{1}{4\pi}\epsilon^{\mu\nu\lambda}A_{\mu}\partial_{\nu}A_{\lambda}$, so the system is an insulator with a unique ground state but with an IQHE response. Note that this looks somewhat similar to some of the discussion of [14], which sees a transition from a $\nu=\frac{2}{3}$ fermionic FCI either to a CDW insulator with no Hall response or one with an IQHE response. We leave more study of this phase to further work.
We should note that Barkeshli and McGreevy [11] similarly discussed a phase “in between” FCI and SF, theirs with (non-BdG) parton Chern numbers $(\mathcal{C}^{\text{P}}_{1},\mathcal{C}^{\text{P}}_{2})=(1,0)$ which turns out to be a CDW Mott insulator as discussed above — one has to break translation symmetry to have the Chern number jump by one unit. Our phase is fundamentally different since it requires parton pairing, which this Mott insulator does not.
IV.4 BdG many-body Chern number
@@
$\left.\middle|\Omega\right\rangle=\prod_{\mathbf{k}}^{\prime}\det U_{\mathbf{k}}e^{(g_{\mathbf{k}})_{\sigma\rho}f_{\sigma,\mathbf{k}}^{\dagger}f^{\dagger}_{\rho,\mathbf{-k}}}\left.\middle|\emptyset\right\rangle.$ (S14)
@@
V Self-consistent mean-field
@@
We see four points where the Berry curvature is concentrated, corresponding to the four band gap closures. Denoting the momentum of one by $\mathbf{k}$, the BdG nature of the Hamiltonian leads to a closure at $-\mathbf{k}$. Furthermore, as discussed above, projective translational symmetry requires gap closures at $\mathbf{k}+(0,\pi)$ and $-\mathbf{k}+(0,\pi)$. The closures seen in Fig S3 are consistent with this, confirming that the transition seen is of this projective-translation-protected gap closure type. A similar construction of four gap closures was found in a related scenario in [13].

### v2

@@ V Comparison to DMRG on the hexagonal model
VI Paired state details
VI.1 Bogoliubov transformation and the mean-field wavefunction form
@@
where $\Psi_{\mathbf{k}}=(f_{1,\mathbf{k}},\ldots,f_{D,\mathbf{k}},f^{\dagger}_{1,\mathbf{-k}},\ldots,f^{\dagger}_{D,-\mathbf{k}})$ and $\Delta(-\mathbf{k})=-\Delta^{T}(\mathbf{k})$. Both $T({\bf k})$ and $\Delta({\bf k})$ are $D\times D$ matrices in this sublattice space. So the matrix in Eq.S10 is $2D\times 2D$ so we get $2D$ BdG bands. From the BdG form, we expect the MF GS $\left.\middle|\Omega\right\rangle$ to fill $D$ of these bands. What follows is a generalization of Read and Green's logic [24] to multiple lattice sites. Following Read and Green, the $D$ empty bands correspond to operators
@@
for l=1Dl=1...D. The Latin indices represent bands and the Greek indices represent (parton) lattice sites in the model. The ground state is then determined by the constraint that $\alpha_{l\mathbf{k}}\left.\middle|\Omega\right\rangle=0$ for all $\mathbf{k}$ and $l$. It may be written as (prime on the product indicates to only multiply over one of each $\mathbf{k},-\mathbf{k}$ pair)
[…]
@@
[…]
Then, writing $G=(g_{\mathbf{k}})_{\sigma\rho}f_{\sigma,\mathbf{k}}^{\dagger}f^{\dagger}_{\rho,\mathbf{{-k}}}$, we can see that $f_{\beta,\mathbf{k}},G]=(g_{\mathbf{k}})_{\beta\rho}f^{\dagger}_{\rho,\mathbf{-k}}$ from which it follows (with $e^{-A}Be^{A}=B+[B,A]+\frac{1}{2!}[[B,A],A]+\frac{1}{3!}[[[B,A],A],A]+\ldots$) that $f_{\beta,\mathbf{k}}e^{G}=e^{G}\left(f_{\beta,\mathbf{k}}+(g_{\mathbf{k}})_{\beta\rho}f^{\dagger}_{\rho,\mathbf{-k}}\right)$.Using that in Eq. S14 along with $f_{\beta,\mathbf{k}}\left.\middle|\emptyset\right\rangle=0$ and $e^{G}f^{\dagger}_{\rho,-\mathbf{k}}=f^{\dagger}_{\rho,-\mathbf{k}}e^{G}$, we get that
@@
VI.2 Recovering mean-field expectations
@@
This is equivalent to the more well-known condition requiring that the Bogoliubovs must obey both $\{\alpha_{l\mathbf{k}},\alpha^{\dagger}_{m\mathbf{q}}\}=\delta_{lm}\delta_{\mathbf{kq}}$ and $\{\alpha_{l\mathbf{k}},\alpha_{m\mathbf{q}}\}=0$. In what follows, we interpret the $2D\times D$ matrix $\begin{pmatrix}U_{\mathbf{k}}^{\dagger}\\ -V_{\mathbf{k}}^{\dagger}\end{pmatrix}$ and similar constructions as a set of $D$ vectors with $2D$ components each. Then, the condition that $U_{\mathbf{k}}g_{\mathbf{k}}=V_{\mathbf{k}}$, is really just the requirement that $\begin{pmatrix}U_{\mathbf{k}}^{\dagger}\\ -V_{\mathbf{k}}^{\dagger}\end{pmatrix}$ is orthogonal to $\begin{pmatrix}g_{\mathbf{k}}\\ \mathbb{1}\end{pmatrix}$, ie that each of the $D$ vectors in one is orthogonal to each of the $D$ in the other. Because $\begin{pmatrix}g_{\mathbf{k}}\\ \mathbb{1}\end{pmatrix}$ is of rank $D$ (guaranteed by the $D\times D$ identity), the dimension of the space orthogonal to it has dimension 2DD=D2D-D=D. In fact, it may be checked that a set of $D$ linearly independent vectors orthogonal to it are just the columns of $\begin{pmatrix}\mathbb{1}\\ -g_{\mathbf{k}}^{\dagger}\end{pmatrix}$. So all the vectors of $\begin{pmatrix}U_{\mathbf{k}}^{\dagger}\\ -V_{\mathbf{k}}^{\dagger}\end{pmatrix}$ are within the span of the vectors in $\begin{pmatrix}\mathbb{1}\\ -g_{\mathbf{k}}^{\dagger}\end{pmatrix}$. Furthermore, unitarity of Eq. S17 necessitates among other things $U_{\mathbf{k}}^{\dagger}U_{\mathbf{k}}+V_{\mathbf{k}}^{\dagger}V_{\mathbf{k}}=\mathbb{1}$, in other words the vectors in $\begin{pmatrix}U_{\mathbf{k}}^{\dagger}\\ -V_{\mathbf{k}}^{\dagger}\end{pmatrix}$ are orthonormal. This all implies that we may take
@@
So given a $g(\mathbf{r})$, we have an algorithm which both figures out the appropriate normalization and then computes the BdG Chern number, pairing strength and can also in principle compute any other correlation function at the mean-field level, as Wick's theorem holds there. Still, we might want to further address the question of why one singles out the particular mean field state where $N(\theta^{*})=NK$. After all, the other $\theta$ which represent different parton mean-field states produce the same projected bosonic wavefunction. The main reason is that the mean-field parton wavefunction lies in a larger Hilbert space, much of which is unphysical — demanding that $N(\theta^{*})=NK$ guarantees that as large a component of that wavefunction is actually in the physical subspace as possible, meaning that mean-field values evaluated at $\theta^{*}$ should be more meaningful for the properties of the physical state — we again note that one has to be careful when interpreting mean-field values physically.
VI.3 Paired Chern numbers and transitions
In the unpaired, $\Delta=0$ and uncorrelated limit, the BdG Chern number is just $\mathcal{C}^{\text{Bdg}}=2(\mathcal{C}^{\text{P}}_{1}+\mathcal{C}^{\text{P}}_{2})$. To see why the factor of $2$, we note that the BdG Chern number counts the number of Majorana edge modes, while the usual Chern numbers count the Dirac edge modes. So we expect $\mathcal{C}^{\text{Bdg}}_{\text{SF}}=0$ and $\mathcal{C}^{\text{Bdg}}_{\text{FCI}}=4$ in the unpaired limit. Assuming a small pairing $\Delta$ does not close any gaps, this should also hold at mean-field for the paired states. We still have the projective translation symmetry for partons $T_{1}^{\text{P}}T_{2}^{\text{P}}=-T_{2}^{\text{P}}T_{1}^{\text{P}}$ which means a doubling in the spectrum, implying that the Chern number jumps by two units, or really double the jump that would happen otherwise. Furthermore, the BdG nature of the problem dictates that a band gap closure at a point $\mathbf{k}$ necessitates another closure at $\mathbf{k}$. Following logic similar to [2], we may then argue that a generic band-gap closure occurs at four points in the BZ, $\mathbf{k},\mathbf{k}+(0,\pi),\mathbf{-k}$ and $\mathbf{-k}+(0,\pi)$. This leads to the conclusion that the jump from $\mathcal{C}^{\text{Bdg}}=4$ to $\mathcal{C}^{\text{Bdg}}=0$ is natural and that the SF-FCI transition may occur without fine-tuning.
@@ VI.3 Paired Chern numbers and transitions
One way this $\mathcal{C}^{\text{BdG}}=2$ phase could be realized is with parton $f_{2}$ in a $\mathcal{C}=1$ (non-BdG) band, while $f_{1}$ transitions from a $\mathcal{C}^{\text{P}}=1$ (non-BdG) band into a $\mathcal{C}=0$ (BdG) band via a double-Majorana gap closure as discussed above. Following Barkeshli and McGreevy [3], we characterize this phase by its electromagnetic response. The parton $f_{1}$ feels a gauge field $\frac{1}{2}A_{\mu}+a_{\mu}$ while $f_{2}$ feels $\frac{1}{2}A_{\mu}-a_{\mu}$ (here $A_{\mu}$ is external electromag field and $a_{\mu}$ is the emergent $U(1)$ ``gluing'' gauge field from the parton construction). Since $f_{1}$ is in a paired state, the gauge field it feels, $b_{\mu}=\frac{1}{2}A_{\mu}+a_{\mu}$ gets Higgsed, ie we get a term $cb_{\mu}b^{\mu}\subset\mathcal{L}$. Integrating out the other parton, $f_{2}$ gives the simple IQHE response $\frac{1}{4\pi}\epsilon^{\mu\nu\lambda}(\frac{1}{2}A-a)_{\mu}\partial_{\nu}(\frac{1}{2}A-a)_{\lambda}\subset\mathcal{L}$. Now when integrating out $a_{\mu}$, the most relevant at large distance is the Higgs term which simply sets $b_{\mu}\rightarrow 0$ meaning that we get $a_{\mu}\rightarrow-\frac{1}{2}A_{\mu}$ and so the final electromagnetic response is simply $\mathcal{L}=\frac{1}{4\pi}\epsilon^{\mu\nu\lambda}A_{\mu}\partial_{\nu}A_{\lambda}$, so the system is an insulator with a unique ground state but with an IQHE response. Note that this looks somewhat similar to some of the discussion of [28], which sees a transition from a $\nu=\frac{2}{3}$ fermionic FCI either to a CDW insulator with no Hall response or one with an IQHE response. We leave more study of this phase to further work.
We should note that Barkeshli and McGreevy [3] similarly discussed a phase ``in between'' FCI and SF, theirs with (non-BdG) parton Chern numbers $(\mathcal{C}^{\text{P}}_{1},\mathcal{C}^{\text{P}}_{2})=(1,0)$ which turns out to be a CDW Mott insulator as discussed above — one has to break translation symmetry to have the Chern number jump by one unit. Our phase is fundamentally different since it requires parton pairing, which this Mott insulator does not.
VI.4 BdG many-body Chern number
@@
[…]
@@
VII Self-consistent mean-field
@@
We see four points where the Berry curvature is concentrated, corresponding to the four band gap closures. Denoting the momentum of one by $\mathbf{k}$, the BdG nature of the Hamiltonian leads to a closure at $\mathbf{k}$. Furthermore, as discussed above, projective translational symmetry requires gap closures at $\mathbf{k}+(0,\pi)$ and $\mathbf{k}+(0,\pi)$. The closures seen in Fig S5 are consistent with this, confirming that the transition seen is of this projective-translation-protected gap closure type. A similar construction of four gap closures was found in a related scenario in [2].